OpenStudy (anonymous):

determine the domain of each of the following. F(x)=4x+4 F(x)=x-3/x^2-16 F(x)=squ3x-6

OpenStudy (anonymous):

Help

OpenStudy (cwrw238):

first : - all real values of x second:- all real values except x = -4 and x = 4 third:- is that the square of (3x - 6) ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

and yes it is

OpenStudy (anonymous):

what would be the 3rd one

OpenStudy (cwrw238):

i think its all real x except x = -2 because that would make 3x-6 = 0

OpenStudy (cwrw238):

- ok i just checked - square of 0 is 0 so domain of third function is all real x

OpenStudy (anonymous):

thanks could you answer these Evaluate the following given f(x) = 4x + 3 and g(x) = x2 + 4x – 1. Determine the domain of each. 9. f(4) 10. g(2) 11. f(x + 3) 12. g(x + 2) 13. (f + g)(x) 14. (f – g)(x) 15. (fg)(x) 16. (f/g)(x)